both sides of a parallelogram measure 8 cm and 10 cm between them and form an angles of 120 degrees. calculate the length of their diagonal
Answer by scott8148(5880)

let x=diagonal __ using the law of cosines __ x^2=8^2+10^2-(2*8*10*cos(120°))

x^2=64+100-(160(-1/2)) __ x^2=244 __ x=sqrt(244) __ x=2sqrt(61)